A particular brand of diet margarine was analyzed to determine the level of polyunsaturated


fatty acid (in percent). A sample of 6 packages resulted in the following data: 16.8, 17.2, 17.4,


16.9, 16.5, 17.1.Test the hypothesis Ho: µ=17 against H1: µ ≠ 17 using α=0.01.

1
Expert's answer
2022-01-19T17:47:37-0500

Solution:

The parameter of interest is the true mean level of polyunsaturated fatty acid,"\mu" .

"\mathrm{H}_0: \mu=17

\\ \mathrm{H}_1: \mu \neq 17

\\ \alpha=0.01

\\ t_{0}=\dfrac{\bar{x}-\mu_{0}}{\sqrt{n}}"

Reject "H_{0}" if "t_{0}<-t_{\alpha / 2,n-1}" where "-t_{0.005,5}=-4.032" or "t_{0}>t_{0.005,5}=4.032"

"\bar{x}=16.98, \mathrm{~s}=0.318,\, \mathrm{n}=6 \\ \quad t_{0}=\dfrac{16.98-17}{0.318 / \sqrt{6}}=-0.154"

Since "-4.032<-0.154<4.032" , do not reject the null hypothesis and conclude the true mean level is not significantly different from 17 % at "\alpha=0.01" .